OpenStudy (anonymous):

Simplify your answer should only contain only positive exponents 4n times 4n^-3 times 4n^0

OpenStudy (anonymous):

\[ 4n \cdot 4n^{-3} \cdot 4n^0 = 4n \cdot \bigg(4 \cdot \frac{1}{n^3} \bigg) \cdot (4 \cdot 1) = \frac{4^3}{n^2} = \frac{64}{n^2} \]